OpenStudy (anonymous):

What is the solution to the equation 9^(x – 1) = 27 ? What is the solution to the equation 9^(x – 1) = 27 ? @Mathematics

OpenStudy (anonymous):

I got x = 2.5

OpenStudy (lilg132):

yes x = 2.5

OpenStudy (anonymous):

3^2(x-1) = 3^3 2x-2 = 3 2x=5 x=5/2

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

yw